• Population A Age (years) Population Number of deaths 15-29 1,000 34 30-49 4,000 26 50+ 6,000 121 • Population B Age (years) Population Number of deaths 15-29 5,000 120 30-49 2,000 10 50+ 1,000 10 5. Calculate the age-group specific rate ratios for population A compared with population B.
Answer #1

Population A age 15-29 years number of deaths - 100034

Population A age 30 -49 years number of deaths - 400026

Population A age 50+ years number of deaths - 6000121

Population B age 15-29 years number of deaths - 5000120

Population B age 30-49 years number of deaths - 200010

Population B age 50+ years number of deaths - 100010

Calculate the age-specific (mortality) rates for each age group by dividing the number of deaths by the respective population.

Total number of population A - 6500181

Now , for age 15 - 29 years the age specific mortality ration is = 100034 / 6500181 = 0.015

For age 30 -49 years the age specific mortality rate ratio = 400026/ 6500181 = 0.0615

For age 50+ years ,the age specific mortality rate ratio = 6000121 / 6500181= 0.923

Total number of population B = 5300140

For age 15-29 years specific mortality rate ratio is - 5000120/ 5300140 = 0.94

For age 30-49 years specific mortality rate ratio - 200010/ 5300140 = 0.037

For age 50+ years specific mortality rate ratio - 100010 / 5300140 = 0.018